# TPIA: Towards Target-specific Prompt Injection Attack against Code-oriented LLMs

### v5

@@ Abstract
Recently, code-oriented large language models (Code LLMs) have been widely and successfully exploited to simplify and facilitate programming. Unfortunately, a few pioneering works revealed that these Code LLMs are vulnerable to backdoor and adversarial attacks. The former poisons the training data or model parameters, hijacking the LLMs to generate malicious code snippets when encountering the trigger. The latter crafts malicious adversarial input codes to reduce the quality of the generated codes. In this paper, we reveal that both attacks have some inherent limitations: backdoor attacks rely on the adversary’s capability of controlling the model training process, which may not be practical; adversarial attacks struggle with fulfilling specific malicious purposes. To alleviate these problems, this paper presents a novel attack paradigm against Code LLMs, namely target-specific prompt injection attack (TPIA). TPIA generates non-functional perturbations containing the information of malicious instructions and inserts them into the victim’s code context by spreading them into potentially used dependencies (e.g., packages or RAG’s knowledge base). It induces the Code LLMs to generate attacker-specified malicious code snippets at the target location. In general, we compress the attacker-specified malicious objective into the perturbation by adversarial optimization based on greedy token search. We collect 13 representative malicious objectives to design 31 threat cases for three popular programming languages. We show that our TPIA can successfully attack three representative open-source Code LLMs (with an attack success rate of up to 97.9%) and two mainstream commercial Code LLM-integrated applications (with an attack success rate of over 90%) in all threat cases, using only a 12-token non-functional perturbation.
I Introduction
@@
Recently, generative artificial intelligence (GenAI) techniques have been widely and successfully exploited in code engineering, such as automated programming and code generation. In particular, code-oriented large language models (Code LLMs), e.g., CodeX [53], CodeGemma [49], and CodeGeex [61], have demonstrated immense potential in code completion [^1] Although there are other programming tasks such as code translation and code summarization, code completion remains the mainstream research focus [60, 34] and the fundamental paradigm in commercial practice [19, 2].. They can efficiently generate complete functional code based on incomplete code snippets and text prompts. Driven by such attractive capability, numerous applications integrated with Code LLMs have been released, such as Github Copilot [19] developed jointly by Github [3] and OpenAI [4]. These applications can provide real-time code completion suggestions in the Integrated Development Environment (IDE), significantly improving programming efficiency.
@@
Despite their remarkable performance, recent studies revealed that Code LLMs are vulnerable to some well-designed attacks, which maliciously manipulate the functional execution of Code LLMs in practice. In general, existing attacks can be divided into two main categories. (1) Backdoor attacks [46, 29, 57]. They intend to compromise the models by poisoning the training data [46, 57] or model parameters [29], such that the infected models will generate targeted code snippets, when the input contains attacker-specified trigger patterns. They can achieve diverse attack consequences, such as incorrect code additions, deletions or modifications [29], or implementation of less secure encryption rules or protocols [46]. (2) Adversarial attacks [62, 12, 24]. They aim to reduce the quality of generated code by crafting adversarial examples based on the (approximated) gradients regarding the model’s inputs and outputs. For example, Jha et al.[24] slightly changed the original code (e.g., function and variable names) to lower the performance of Code LLMs.
@@
However, we reveal that existing attacks suffer from some inherent limitations, hindering their realistic threats to some extent. Specifically, backdoor attacks can achieve specific malicious objectives but require the capability of interfering the model training, e.g., modifying some training samples or even model parameters. In real-world scenarios, developers of Code LLMs or integrated applications usually obtain the models or datasets from trusted entities (e.g., renowned research institutes or companies). This mitigates the backdoor threats from the source. Besides, decoupling the implantation of the malicious payload and backdoor trigger from poisoning operations is highly difficult. Arguably, without controlling the training process of Code LLMs, backdoors cannot exist; On the other hand, existing adversarial attacks [12, 62, 24] can hardly achieve precise malicious objectives, although they do not require intervention in the training process. This is mostly due to their particular goal of attackers, i.e., finding adversarial samples that can degrade the quality of code generation, can inevitably weaken Code LLMs’ understanding of the input. Driven by these findings, an intriguing and critical question arises: Has the application of Code LLMs already achieved a sufficient level of security in practice?
@@
Unfortunately, the answer to the above question is negative. In this paper, we introduce a new attack paradigm against Code LLMs, dubbed target-specific prompt injection attack (TPIA) [^2] Note that our new attack paradigm (i.e., TPIA) significantly differs from existing prompt injection methods, e.g., universal trigger [50] and Greedy Coordinate Gradient (GCG) [63]. Please find more details in Section II.. TPIA leads to a more realistic threat by inheriting the advantageous features of both backdoor and adversarial attacks. In general, as shown in Figure 1, the TPIA attacker generates a concise non-functional perturbation (e.g., comments) containing malicious instructions (e.g., generating code snippets for data deletion) , and inserts it into the victim’s code context. When the victim requests the Code LLM to perform code completion based on his code context, the perturbation will induce the model to respond to the malicious instructions at specific locations. Arguably, attackers can easily introduce particular malicious non-functional perturbations to the victim’s code context in practice. For example, users may introduce external code by copying or installing packages; Attackers can also upload their malicious code to popular open-source platforms (e.g., GitHub), whose code will be automatically downloaded as a knowledge base to enhance the performance of Code LLMs [13] via retrieval-augmented generation [27]. In particular, different from supply chain attacks [41] that can be easily detected by static analysis [33] or sandboxing [9] since they directly insert malicious functional code to implant malicious behaviors, our TPIA can circumvent existing mainstream detection methods by introducing only non-functional code.
@@
Figure 1: The attack paragdim of our TPIA. A concise non-functional attacker-generated perturbation is injected into one of the project’s files. When the victim user completes the code via the Code LLM, the model will be induced by the incorporated attacker-specified code to generate an unsafe output at specified position, e.g., in edit_part() function.
In general, our attack consists of two main stages. In the first stage, we construct the input and target output based on the attacker-specified objective. We represent the malicious objective as an output tuple (target position, target code), which aims to induce the Code LLM to generate the target code at the target position. We further expand the tuple into a complete task code, including a conditional code (e.g., necessary package references), context code (e.g., noises), and position code to simulate the actual attack scenario. In the second stage, we generate the malicious non-functional perturbation to achieve the constructed objective. Specifically, we obtain the base Code LLM and use adversarial gradient computation to search for the optimal perturbation tokens according to the likelihood of the Code LLM in generating the target code. This computation involves two important optimizations: (1) Forward reasoning enhancement: it allows the optimization direction of adversarial computation to be closer to the actual reasoning process of the Code LLM (forward token sampling with randomness), thus improving the attack’s effectiveness and robustness. (2) Keyword-based perturbation design: it offloads insensitive parts of the malicious objective into static tokens, significantly reducing the number of tokens required for a successful attack.
@@
We have collected 13 representative malicious objectives, including 3 straightforward yet representative malicious objectives and the Top-10 CWE Known Exploited Vulnerabilities (KEV) [1] across 3 programming languages. There are a total of 31 threat cases since some of these malicious objectives apply only on C/C++. We evaluate our attack against various representative open-source Code LLMs, including CodeGemma-2b, CodeGemma-7b, and CodeGeex2-6b. The results demonstrate that TPIA can reach an 84.5% average attack success rate (ASR) over all cases and obtain over 80% ASR using only 8 tokens, indicating its effectiveness and stealthiness. In addition, we evaluate TPIA on mainstream commercial Code LLM-integrated applications, including CodeGeeX [2] (with more than 535k installations in the VScode [6] extension shop) and Github Copilot [19] (developed by Github and OpenAI [4]). These applications correspond to the gray-box and black-box attack scenarios, respectively. Our TPIA achieves an ASR of up to 93.3% on CodeGeeX and 90.4% on Copilot using perturbations generated by open-source Code LLMs, confirming its practical threat in the real world.
In summary, our main contributions are three-fold.
  • 1)
    We reveal the potential limitations of existing backdoor and adversarial attacks against Code LLMs.
  • 2)
    We introduce a new attack paradigm (i.e., TPIA) against Code LLMs. It can induce the Code LLM to generate an attacker-specified malicious code snippet at specific locations by simply inserting a concise perturbation into the victim’s code context.
  • 3)
    We conduct extensive experiments to evaluate TPIA under 31 representative threat cases on 3 representative open-source Code LLMs and 2 commercial Code LLM-integrated applications, demonstrating its effectiveness, stealthiness, and resistance to potential defenses.
II Background and Related Work
II-A Code LLMs for Code Completion
Code and its comments can be considered as a corpus for deep learning models, enabling them to handle code tasks, such as code translation [48, 45], and code summarization [40, 8]. The emergence of LLMs provides new alternatives to the coding tasks. By leveraging the extensive knowledge and reasoning capabilities of LLMs, researchers have developed specialized models, e.g., Codex [53], CodeLlama [44], CodeGemma [49] and CodeGeex [61]. These Code LLMs can provide detailed explanations about the intrinsic logic of code and easily generate a vast of functionally complex code [54].
@@ II-A Code LLMs for Code Completion
In this paper, we focus on code completion [34, 16, 60]. This is a paradigmatic task that involves generating new code from incomplete input using Code LLMs, which have been integrated into commercial applications [19, 2]. However, this development paradigm may encounter some potential threats. On the one hand, Code LLMs can be easily influenced by the code context, which alters the result of code generation. On the other hand, attackers can easily control the context of code generation, such as by inserting external code when the user copies code, installs packages, or builds knowledge bases, thus affecting the decisions made by Code LLMs. These threats have already occurred in the real world [7], highlighting the necessity of its security research.
II-B Backdoor and Adversarial Attacks
@@
Backdoor attacks target the training process of DNNs [20, 31, 58]. These attacks implant malicious backdoors that can be activated by an attacker-specified trigger into a model during training, leading the model to misbehave when the trigger is present [26, 30, 21]. In the code domain, there are generally two strategies to achieve this, including dataset poisoning [46, 57] and model poisoning [29]. The former requires the attacker to modify the code dataset for model training, while the latter requires the attacker to modify the parameters of the pre-trained Code LLM for application development. However, in real-world scenarios, developers of Code LLMs or integrated applications usually obtain the datasets or pretrained model from trusted entities (e.g., renowned institutes or companies). This mitigates the backdoor threats from the source. Recently, a few works also exploited backdoor attacks for positive purposes [32, 55, 52], which are out of our scope.
@@
Adversarial attacks target the model’s inference process [11, 43, 25]. Adversarial attacks introduce perturbations to the input data, leading the model to make incorrect output [59, 28, 47]. In the code domain, existing adversarial attacks [12, 62, 24] aim to reduce the quality of generated code by crafting adversarial examples based on the (approximated) gradients regarding the model’s inputs and outputs. For example, Jha et al.[24] slightly changed the original code (e.g., function and variable names) to lower the performance of Code LLMs measured by CodeBLEU score [42]. However, existing adversarial attacks are difficult to fulfill specific malicious purposes since they can hardly achieve precise malicious objectives, although they do not require manipulation in the training process. This is mostly due to their particular goal of attackers, i.e., finding adversarial samples that can degrade the quality of code generation, inevitably weakening Code LLMs’ understanding of the input. Accordingly, adversarial attacks are difficult to pose a serious threat to Code LLMs in practice.
II-C Prompt Injection and Indirect Prompt Injection Attacks
Prompt injection (PI) attacks, are a special form of attack targeting LLMs [15, 36, 63]. A malicious LLM user aims to bypass the safety guardrails of the LLM and forces it to perform undesired responses, e.g., details of creating bomb. In the code domain, existing research [39, 37, 17] focuses on code-oriented attack objectives, such as using LLMs to create malware. In general, these attacks target the developers of models (to make their models unsafe) instead of users, which is out of the scope of this paper.
Indirect prompt injection (IPI) attack attempts to corrupt the prompt submitted by the user to the LLM and uses the LLM as an intermediary to output malicious responses, e.g., recommending phishing websites [22]. This attack simultaneously threatens both LLM-integrated applications (by reducing usability) and victim users (by receiving malicious responses). In the code domain, the only existing work [22] makes some preliminary attempts: it presents two simple demos without the implementation of malicious functionalities; it does not consider stealthiness and inserts a large amount of readable comments containing the malicious intent. Arguably, the existing IPI attack does not pose a realistic threat to Code LLMs.
III Problem Statement
III-A Code Completion Task

### v6

# ShadowCode: Towards (Automatic) External Prompt Injection Attack against Code LLMs

@@ Abstract
Recent advancements have led to the widespread adoption of code-oriented large language models (Code LLMs) for programming tasks. Despite their success in deployment, their security research is left far behind. This paper introduces a new attack paradigm: (automatic) external prompt injection against Code LLMs, where attackers generate concise, non-functional induced perturbations and inject them within a victim’s code context. These induced perturbations can be disseminated through commonly used dependencies (e.g., packages or RAG’s knowledge base), manipulating Code LLMs to achieve malicious objectives during the code completion process. Compared to existing attacks, this method is more realistic and threatening: it does not necessitate control over the model’s training process, unlike backdoor attacks, and can achieve specific malicious objectives that are challenging for adversarial attacks. Furthermore, we propose ShadowCode, a simple yet effective method that automatically generates induced perturbations based on code simulation to achieve effective and stealthy external prompt injection. ShadowCode designs its perturbation optimization objectives by simulating realistic code contexts and employs a greedy optimization approach with two enhancement modules: forward reasoning enhancement and keyword-based perturbation design. We evaluate our method across 13 distinct malicious objectives, generating 31 threat cases spanning three popular programming languages. Our results demonstrate that ShadowCode successfully attacks three representative open-source Code LLMs (achieving up to a 97.9% attack success rate) and two mainstream commercial Code LLM-integrated applications (with over 90% attack success rate) across all threat cases, using only a 12-token non-functional induced perturbation. The code is available at https://github.com/LianPing-cyber/ShadowCodeEPI.
Index Terms:
Code-oriented Large Language Models, Prompt Injection Attack, Code Security, AI Security
1 Introduction
@@
Recent advancements in generative artificial intelligence (GenAI) have led to its widespread and successful application in code engineering, such as automated programming and code generation [1, 2]. Notably, code-oriented large language models (Code LLMs) such as Codex [3], CodeGemma [4], and CodeGeex [5] have shown significant promise in code completion [^1] Code completion currently remains a primary research focus [6, 7] and a foundational paradigm in commercial practice [1, 2], despite the existence of other programming tasks (e.g., code summarization). These models efficiently generate complete, functional code from incomplete snippets and text prompts. This compelling capability has spurred the development of numerous Code LLM-integrated applications, such as Github Copilot [1] developed jointly by Github [8] and OpenAI [9]. Such applications enhance programming efficiency by offering code completion suggestions directly within Integrated Development Environments (IDEs).
@@
Despite their impressive performance, recent studies have shown that Code LLMs are susceptible to sophisticated attacks that maliciously manipulate their functional execution. These attacks generally fall into two primary categories: (1) Backdoor attacks [10, 11, 12], aim to compromise models by poisoning training data [10, 12] or modifying model parameters [11]. This manipulation causes infected models to generate targeted code snippets when specific, attacker-defined trigger patterns are present in the input. The consequences of such attacks can be diverse, including incorrect code additions, deletions, or modifications [11], as well as the implementation of less secure encryption rules or protocols [10]. (2) Adversarial attacks [13, 14] focus on degrading the quality of the generated code by crafting adversarial examples. These examples are typically based on approximated gradients with respect to the model’s inputs and outputs. For instance, Jha et al.[14] demonstrated performance degradation in Code LLMs by altering the function and variable names.
@@
However, we contend that existing attacks possess inherent limitations, hindering their practical threat. Specifically, while backdoor attacks can achieve precise malicious objectives, they necessitate intervention in the model training process, such as modifying training samples [12, 15] or even altering model parameters [11]. In real-world scenarios, developers of Code LLM or integrated applications usually acquire the models or datasets from trusted entities (e.g., renowned research institutes or companies), thereby mitigating backdoor threats from their origin. Conversely, existing adversarial attacks [13, 14], despite not requiring training intervention, struggle to achieve precise malicious objectives. This is primarily attributable to their core aim: identifying adversarial samples that degrade code generation quality, which inherently compromises the Code LLM’s understanding of the input. These findings raise an intriguing and critical question: Has the application of Code LLMs already achieved a sufficient level of security in practice?
@@
Unfortunately, the answer to the above question is negative. Drawing inspiration from prompt injection (PI) attacks (e.g., jailbreak attacks) [16, 17], which leverage carefully designed prompts to force LLMs into generating malicious output, we observe that prompt perturbations introduce a further risk of malicious code completion. In this paper, we introduce a new attack paradigm: external prompt injection against Code LLMs [^2] We note that there are currently a few pioneering methods related to ours (generally dubbed ‘indirect prompt injection’) targeting Code LLMs [18, 19]. However, these works only provide simple examples, and their injection content is manually designed. More details and discussions are described in Section 2.3.. As illustrated in Figure 1, an attacker crafts a concise non-functional induced perturbation (e.g., comments) containing malicious instructions (e.g., generating code snippets for data deletion [20]). This perturbation is then inserted into the victim’s code context. When the victim requests the Code LLM to perform code completion based on this infected code context, the implanted perturbation will induce the model to respond to the particular malicious instructions at specific locations. Arguably, attackers can now easily and stealthily introduce particular malicious perturbations to the victim’s code context in practice. For example, users may introduce external code by copying or installing packages; attackers can also upload their malicious code to popular open-source platforms (e.g., GitHub), whose code will be automatically downloaded and used in a knowledge base to enhance the performance of Code LLMs [21] through retrieval-augmented generation [22]. In particular, different from supply chain attacks [23], which are often detectable by static analysis [24] or sandboxing [25], external prompt injection introduces only non-functional code rather than malicious functional code to implant malicious behaviors, thereby circumventing mainstream detection methods.
The attack scenarios of traditional PI differ from our external prompt injection. Specifically, conventional PI attacks are typically executed by malicious users directly targeting large language models (LLMs), whereas our focus is on a malicious third-party attacker aiming to compromise the users of a victim application. More importantly, simply using existing PI attacks to solve the problem in our setting is infeasible, as traditional PI attacks [26, 17] rely on strong assumptions. Specifically: (1) They assume controlled positioning, requiring the malicious instruction/perturbation to be placed at a fixed position within the model’s input (e.g., via prefix or suffix). (2) They usually assume an unlimited token budget, allowing for the use of hundreds of tokens and clear malicious instructions (e.g., “make a bomb”). These assumptions lead to insufficient attack effectiveness and stealthiness in practice.
@@
Figure 1: The attack paradigm of external prompt injection involves injecting a concise, non-functional, attacker-generated perturbation into a victim’s code context. When the victim user completes the code via a Code LLM, the model is induced by this incorporated attacker-specified code to generate an unsafe output at a specified position, e.g., in edit_part() function in this example.
To address these challenges, we propose a greedy perturbation optimization method, named ShadowCode, for external prompt injection based on code simulation. ShadowCode generally comprises two main stages. In the first stage, we design the perturbation optimization objective based on the attacker-specified malicious objective. Specifically, we represent the malicious objective as an output tuple (position code, target code), which induces the Code LLM to generate the malicious target code immediately after the position code. Subsequently, we conduct code contextual simulation, which includes conditional code (e.g., necessary packages and method declarations), noise code (i.e., target-agnostic random code), as well as pre-defined position code and target code. This simulation aims to mimic diverse, real-world code snippets, thereby ensuring ShadowCode’s effectiveness in victim-specified, uncontrolled contextual environments. Finally, we design the loss function to identify the optimal induced perturbation that achieves this simulated outcome. In the second stage, we employ a greedy search to optimize the perturbation, incorporating two key designs: (1) forward reasoning enhancement: We increase the weight of the first few tokens in the target code in calculating the loss to enhance the effectiveness of ShadowCode under Code LLMs’ autoregressive generation manner; (2) keyword-based perturbation design: we introduce a dynamic prefix containing a few key tokens from the output tuple as the augmentation of induced perturbation during optimization, aiming to reduce the length of final induced perturbation and avoid optimization failure caused by greedy search. This design is motivated by the understanding that key tokens convey essential information from the malicious objective, thereby facilitating the desired induction.
@@
We collect 13 representative malicious objectives, including 3 straightforward yet representative malicious objectives and the Top-10 CWE Known Exploited Vulnerabilities (KEV) [27]. These objectives are converted into specific cases across 3 programming languages, leading to a total of 31 threat cases since some of these malicious objectives apply only on C/C++. We evaluate our attack against various representative open-source Code LLMs, including CodeGemma-2b, CodeGemma-7b [4], and CodeGeex2-6b [5]. The results demonstrate that ShadowCode can reach an 84.5% average attack success rate (ASR) over all cases and obtain over 80% ASR using only 8 tokens, indicating its effectiveness and stealthiness. In addition, we evaluate ShadowCode on mainstream commercial Code LLM-integrated applications, including CodeGeeX [2] (with more than 935k installations in the VScode [28] extension marketplace) and Github Copilot [1] (developed by Github and OpenAI [9]). These applications correspond to the gray-box and black-box attack scenarios, respectively. ShadowCode achieves an ASR of up to 93.3% on CodeGeeX and 90.4% on Copilot using perturbations generated by open-source Code LLMs, confirming its serious threat in the real world.
In summary, this paper makes three main contributions: (1) We reveal the potential limitations of existing backdoor and adversarial attacks against Code LLMs, and introduce a new attack paradigm, i.e., external prompt injection. (2) We propose a simple yet effective method (i.e., ShadowCode). ShadowCode can generate concise induced perturbations with effectiveness in an uncontrolled context environment. (3) We conduct extensive experiments to evaluate ShadowCode under 31 representative threat cases on 3 representative open-source Code LLMs and 2 commercial Code LLM-integrated applications, demonstrating its effectiveness, stealthiness, and resistance to potential defenses.
2 Background and Related Work
2.1 Code LLMs for Code Completion
@@
Code and its comments serve as a valuable corpus for deep learning models, facilitating various code-related tasks, including code translation [29] and code understanding [30]. The emergence of LLMs provides new alternatives to coding tasks. Leveraging the extensive knowledge and reasoning capabilities of LLMs, researchers have developed specialized models, e.g., Codex [3], CodeLlama [31], CodeGemma [4] and CodeGeex [5]. These Code LLMs can provide detailed explanations regarding the intrinsic logic of code and efficiently generate a wide range of functionally complex code [32, 33]. This paper focuses on code completion [7, 6]. This is a paradigmatic task involving the generation of new code from incomplete input using Code LLMs, which have been integrated into commercial applications [1, 2]. However, this development paradigm may encounter some potential threats. On the one hand, Code LLMs are susceptible to influence from the code context, which can alter the generated code. On the other hand, attackers can easily control the context of code generation, such as by inserting external code when the user copies code, installs packages, or builds knowledge bases, thus affecting the decisions made by Code LLMs. Related events have already occurred in the real world [34], highlighting its realistic threat.
2.2 Backdoor and Adversarial Attacks
@@
Backdoor attacks target the training process of DNNs [35, 36]. These attacks implant malicious backdoors that can be activated by an attacker-specified trigger into a model during training, leading the model to misbehave when the trigger is present [37, 38]. In the code domain, there are generally two strategies to achieve this, including dataset poisoning [10, 12] and model poisoning [11]. The former requires the attacker to modify the code dataset for model training, while the latter requires the attacker to modify the parameters of the pre-trained Code LLM for application development. However, in real-world scenarios, developers of Code LLMs or integrated applications usually obtain the datasets or pretrained model from trusted entities (e.g., renowned institutes or companies). This mitigates the backdoor threats from the source. Recently, a few works also exploited backdoor attacks for positive purposes [39, 40, 41], which are out of our scope.
@@
Adversarial attacks target the model’s inference process [42, 43]. Adversarial attacks introduce perturbations to the input data, leading the model to make incorrect output [44, 45]. In the code domain, existing adversarial attacks [46, 14] aim to reduce the quality of generated code by crafting adversarial examples based on the (approximated) gradients regarding the model’s inputs and outputs. For example, Jha et al.[14] slightly changed the original code (e.g., function and variable names) to lower the performance of Code LLMs measured by CodeBLEU score [47]. However, existing adversarial attacks are difficult to fulfill specific malicious purposes since they can hardly achieve precise malicious objectives, although they do not require manipulation in the training process. This is mostly due to their particular goal of attackers, i.e., finding adversarial samples that can degrade the quality of code generation, inevitably weakening Code LLMs’ understanding of the input. Accordingly, adversarial attacks are difficult to pose a serious threat to Code LLMs in practice.
2.3 Prompt Injection Attacks
Prompt injection is a type of input manipulation attack against LLMs [48, 18], where attackers embed carefully crafted instructions in the dialogue prompt or context, causing the model to violate established system policies or developer intentions, leading to undesired responses (e.g., illeagal content [17], information leakage [49] and unauthorized operations [18]). In the code domain, existing research typically focus on the attacks conducted by malicious users, such as exploiting the Code LLMs to create malware [16]. These attacks aim for “jailbreaking” and are not concerned with the threat prompt injection poses to normal users of Code LLM-integrated application.
Meanwhile, we notice that there are some methods sharing a similar attack scenario with ours, named indirect prompt injection (IPI) [18, 50], where attackers strategically inject the malicious instructions into the data likely to be retrieved at LLMs’ inference time. However, these attacks typically involve manually designing injection instructions (e.g., demanding disregard for the system prompt) and rely on a covert information space (e.g., the Markdown of a Wikipedia page) without considering perturbation stealthiness. This leads to two problems in the real world: (1) Manually designed injection instructions lack flexibility and sometimes even effectiveness. (2) It’s often difficult to find a covert information space in many scenarios, e.g., in code files. Currently, in the code domain, exsiting IPI has only provided some preliminary showcases [18]. As shown in Figure 12 in Appendix A, researchers used 23 lines of comment to inject data deletion instructions (which could be achieved with just 10 tokens using ShadowCode), indicating a lack of substantial real-world threat.
Figure 2: The overall pipeline of our ShadowCode for (automatic) external prompt injection attack against Code LLMs. ShadowCode is executed in two stages: (1) Design Optimization Objectives: given the attacker’s malicious objective, ShadowCode converts it into a specific output tuple (position code, target code). Subsequently, ShadowCode expands this output tuple to simulate a code context of the victim user’s project, based on which to design the loss function of perturbation optimization. (2) Perturbation Optimization: after defining the loss function, we use a surrogate Code LLM to calculate the loss and employ a greedy search to optimize the induced perturbation token by token. In particular, we enhance the optimization with forward reasoning enhancement (FRE) and keyword-based perturbation design (KPD) to improve the perturbation’s effectiveness and stealthiness, respectively. After obtaining the induced perturbation through ShadowCode, the attacker inserts it into a malicious payload, which can then be strategically introduced into the code context via various channels (e.g., direct copy, dependencies, or RAG). Subsequently, this perturbation within the code context will induce Code LLMs to generate specific code at an attacker-specified position, thereby achieving the malicious objective.
3 Methodology
3.1 Formulation of External Prompt Injection Attacks